# Discriminative Finetuning of Generative Large Language Models without Reward Models and Preference Data

### v1

@@ Abstract
Supervised fine-tuning (SFT) followed by preference optimization (PO) denoted by SFT→PO has become the standard for improving pretrained large language models (LLMs), with PO demonstrating significant performance gains. However, PO methods rely on either human-labeled preference data or a strong reward model to generate preference data. Can we fine-tune LLMs without preference data or reward models while achieving competitive performance to SFT→PO? We address this question by introducing Discriminative Fine-Tuning (DFT), a novel approach that eliminates the need for preference data. Unlike SFT, which employs a generative approach and overlooks negative data, DFT adopts a discriminative paradigm that that increases the probability of positive answers while suppressing potentially negative ones, shifting from token prediction to data prediction. Our contributions include: (i) a discriminative probabilistic framework for fine-tuning LLMs by explicitly modeling the discriminative likelihood of an answer among all possible outputs given an input; (ii) efficient algorithms to optimize this discriminative likelihood; and (iii) extensive experiments demonstrating DFT’s effectiveness, achieving performance better than SFT and comparable to if not better than SFT→PO. The code can be found at https://github.com/PenGuln/DFT.
@@ 1 Introduction
The current standard for fine-tuning LLMs typically involves supervised fine-tuning (SFT) followed by preference optimization (PO), including techniques such as reinforcement learning with human feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024). In this approach, SFT first aligns the model with supervised data, and PO further refines the model using preference data labeled by humans or a reward model, which simulates human preferences. This two-stage process has achieved significant success, particularly in improving human alignment and response quality. However, PO methods often require extensive human annotations or the construction of robust reward models, both of which are resource-intensive and may limit scalability and applicability in highly specialized areas.
This raises an intriguing question: Can we fine-tune LLMs without preference data or reward models while achieving competitive performance to SFT→PO?
To address this question, we propose Discriminative Fine-Tuning (DFT), a novel alternative to SFT→PO that eliminates the need for preference data and reward models. Unlike SFT, which uses a generative approach and overlooks negative examples, DFT adopts a discriminative notion, explicitly discriminating good from “bad” outputs generated by the base model to be finetuned. We formalize this approach by introducing a discriminative probabilistic framework that models the discriminative likelihood of an answer among all possible outputs for a given input. This stands in stark contrast to SFT, which uses a generative probabilistic framework to model only the generative likelihood of individual tokens in the answers. To implement this framework, we develop efficient algorithms to optimize the discriminative likelihood of good answers, ensuring both scalability and practicality. Due to its strong discriminative capability, DFT delivers competitive or superior performance compared to SFT→PO, while eliminating the need for preference data and reward models.
@@ 2 Related Work
Preference Optimization. Pioneering works proposed RL-based PO methods (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). These methods leverage a separate reward model, trained on human-labeled preference data, and optimize the SFT model against it using policy gradient methods such as PPO (Schulman et al., 2017) and REINFORCE (Williams, 1992). Rafailov et al. (2024) proposed direct preference optimization (DPO), which removes the step of training a reward model and directly optimizes a pairwise loss of the policy model on the preference data. Following DPO, many PO methods have been proposed with different loss functions, including R-DPO (Park et al., 2024), CPO (Xu et al., 2024a), IPO (Azar et al., 2024), SimPO (Meng et al., 2024), KTO (Ethayarajh et al., 2024), to name just a few among others (Zhao et al., 2023; Jung et al., 2024).
Several works (Chen et al., 2024b; Yuan et al., 2023; Song et al., 2024) have considered PO with a list of ranked preference data that may be explicitly labeled with a reward value. Rosset et al. (2024) assumed a general preference function is given that can produce a probability telling one output is preferred over another output given an input. Different from these works, we do not assume any preference data or preference model other than annotated input-output pairs.
Finetuning via Self-play. Training a model on its own self-generated responses has been widely explored in the PO stage. For example, many variants of RLHF (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022; Li et al., 2023; Chan et al., 2024; Ji et al., 2024) use on-policy samples produced by the current policy under optimization. Some studies have exhibited benefits of using self-generated data for PO by reducing the distribution gap between the training data and the current model while fostering exploration of diverse response spaces (Xu et al., 2024b; Tajwar et al., 2024; Tang et al., 2024). Moreover, leveraging synthetic data has proven essential for iterative (online) algorithmic improvement of these methods (Xu et al., 2023b; Guo et al., 2024; Yuan et al., 2024; Chen et al., 2024a; Dong et al., 2024). A more closely related work is SPIN (Wu et al., 2024), which uses a similar preference optimization objective as DPO but with data generated by the model to be finetuned as the losing responses. Although we also use self-generated data from the base model to be finetuned as our negative data, our formulation is derived from a discriminative learning framework, making our approach aided by advanced optimization better than the pairwise loss function used in SPIN and other pairwise preference optimization objectives (cf. Section 6).
@@ 4 DFT: Discriminative Finetuning
To address this issue, the current mainstream approach is to finetune the model further using PO on preference data. If humans label the two answers such that $\mathbf{y}\succ\mathbf{y}^{\prime}$, the model might be able to push the likelihood of $\mathbf{y}^{\prime}$ given $\mathbf{x}$ smaller than that of $\mathbf{y}$ given $\mathbf{x}$. As a result, the good answer $\mathbf{y}$ will be generated more likely than the bad answer $\mathbf{y}^{\prime}$.
@@ 4.1 The Scoring Function
The above framework might remind one of the contrastive representation learning (Radford et al., 2021; Wang et al., 2025). However, we cannot directly borrow the same idea of contrastive representation learning to design the scoring function. In particular, contrastive representation learning uses an encoder network $e(\mathbf{x})$ to induce an embedding of any input text $\mathbf{x}$, and computes the scoring function by using the cosine similarity between $e(\mathbf{x})$ and $e(\mathbf{y})$. However, this representation model $e(\cdot)$ is of no use for generative tasks of LLMs.
@@ 4.2 The Optimization Algorithm
Although our DFT formulation (2) and (3) is nearly identical to that of the traditional discriminative probabilistic approach for classification (e.g., logistic regression), the key challenge lies in solving the optimization problem in (3), particularly in handling the second term of $F(\theta)$, where $\mathcal{Y}$ encompasses all possible texts. We address this challenge by advanced optimization techniques. In particular, we cast the second term into an instance of finite-sum coupled compositional optimization framework (FCCO) (Wang & Yang, 2022). The idea is to write the second term of $F(\theta)$ into the form of $\frac{1}{n}\sum_{i=1}^{n}f(\mathbb{E}_{\zeta}g_{i}(\theta;\zeta))$, where $\zeta$ is some random variable. To this end, we introduce a sampling distribution $P_{i}(\cdot)$, which is specified later. Then we define
@@ 7 Experiments
Setup: We evaluate our proposed DFT framework under two distinct training settings. First, we focus on improving the mathematical reasoning capability of a base LLM by using DFT on the MetaMathQA dataset (Yu et al., 2024), which contains 395K samples generated through bootstrapped mathematical questions with augmented reasoning paths. In this setting, we set $B=4$. Second, we fine-tune a base LLM using DFT on the UltraFeedback (UF) dataset (Cui et al., 2023), comprising 61K samples. UF is originally used as a preference dataset, where each data pairs a winning response and a losing response. For SFT and DFT, we regard the winning responses $\mathbf{y}_{w}$ as the ground-truth and discard all losing responses. In this setting, we set $B=2$ and generate $\mathbf{y}^{\prime}$ by adding an adversarial prompt like “You are an unhelpful assistant.” to the input $\mathbf{x}_{i}$ in a chat template (cf. Appendix B.4) and use it as input to the base LLM for generation. For both settings, we use Mistral-7B-v0.1 as our base model. More details of implementation and hyper-parameter tuning are described in Appendix A.
@@ 7.1 Results
Observation 2: DFT variants consistently outperform PO methods on self-play data. In Table 3, we compare DFT(2) with PO approaches using self-play data that is generated by the base model as negative data, including SPIN, SimPO, KTO, and SimPO-SFT, where the last one just combines the SimPO loss and the SFT loss similar to Xu et al. (2024a). Comparing with SimPO-SFT allows us to verify the advantage of our objective over SimPO loss. For these baselines, we use the same generated $\mathbf{y}^{\prime}$ as in DFT as their negative data and finetune the same base model. The results in Table 3 show that these PO approaches on self-play data can not improve SFT. This is different from the observation in Chen et al. (2024c); Xu et al. (2024a), as their experiments are for finetuning an SFT model. Comparing DFT(2) with SPIN and SimPO can justify the effectiveness of our objectives and the optimization algorithm.
@@
RFT-7B 50.3
Qwen-7B 51.6
Mistral-7B-base (8 shots) 52.2 13.1
WizardMath-7B 54.9 10.7
LLaMA-2-70B 56.8 13.5
@@
Table 2: Comparison between DFT, SFT, and PO methods on self-play data in the second training setting.
@@
Table 3: Comparison between DFT and SFT→PO approaches on preference data in the second training setting.
@@
AlpacaEval2 Results. Finally, we briefly discuss the GPT4-as-a-judge evaluation on instruction following by reporting the length-controlled winning rate (LC) on AlapacaEval2. The comparison between DFT(2) with SFT and PO-based methods on self-play data is shown in Figure 2(a), which demonstrates DFT(2) outperforms these baselines. The comparison between DFT(2) and SFT→PO approaches on preference data is shown in Appendix B.1, which shows that DFT(2) is competitive with some PO approaches using the preference data, such as KTO, but worse than SimPO and DPO. However, DFT(2) have much shorter lengths for the outputs with an average length of 1359. In contrast, KTO, DPO, SimPO have average lengths of 1449, 1477, 1868, respectively. The GPT4-as-a-judge evaluation tends to favor outputs that are longer. Nevertheless, DFT has competitive if not better performance on verifiable instruction following benchmark IFEval (cf. Table 3).
@@ 7.2 Ablation Studies
We present some results to illustrate the advantage of our optimization using moving-average estimators $u$, and the effect of the number of generated samples $B$ in each iteration.
@@ Appendix A Implementation Details
A.1 Hyper-parameters Tuning
@@ A.2 Implemenation of SFT→PO
A.3 Training Costs
We compare the training efficiency of different methods, with results shown in Figure 3.
Figure 3: Training time per-iteration of DFT, DFT2 ($B=2$) and SFT and PO methods for finetuning Mistral-7b-base model on UltraFeedback dataset.

### v2

# Discriminative Finetuning of Generative Large Language Models without Reward Models and Human Preference Data

@@ Abstract
Supervised fine-tuning (SFT) has become a crucial step for aligning pretrained large language models (LLMs) using supervised datasets of input-output pairs. However, despite being supervised, SFT is inherently limited by its generative training objective. To address its limitations, the existing common strategy is to follow SFT with a separate phase of preference optimization (PO), which relies on either human-labeled preference data or a strong reward model to guide the learning process. In this paper, we address the limitations of SFT by exploring one of the most successful techniques in conventional supervised learning: discriminative learning. We introduce Discriminative Fine-Tuning (DFT), an improved variant of SFT, which mitigates the burden of collecting human-labeled preference data or training strong reward models. Unlike SFT that employs a generative approach and overlooks negative data, DFT adopts a discriminative paradigm that increases the probability of positive answers while suppressing potentially negative ones, aiming for data prediction instead of token prediction. Our contributions include: (i) a discriminative probabilistic framework for fine-tuning LLMs by explicitly modeling the discriminative likelihood of an answer among all possible outputs given an input; (ii) efficient algorithms to optimize this discriminative likelihood; and (iii) extensive experiments demonstrating DFT’s effectiveness, achieving performance better than SFT and comparable to if not better than SFT→PO. The code can be found at https://github.com/Optimization-AI/DFT.
@@ 1 Introduction
The current standard for aligning LLMs typically involves supervised fine-tuning (SFT) followed by preference optimization (PO) denoted by SFT→PO, including techniques such as reinforcement learning from human feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024). In this approach, SFT first aligns the model with supervised data, and PO further refines the model using preference data labeled by humans or a reward model that simulates human preferences. This two-stage process has achieved significant success, particularly in improving human alignment and response quality. However, PO methods often require extensive human annotations or the construction of robust reward models, both of which are resource-intensive and may limit scalability and applicability in highly specialized areas.
This raises an intriguing question: Can we align LLMs without human preference data or reward models while achieving competitive performance to SFT→PO?
To address this question, we propose Discriminative Fine-Tuning (DFT), a novel alternative to SFT→PO that mitigates the burden of collecting human-labeled preference data or training strong reward models. Unlike SFT, which uses a generative approach and overlooks negative examples, DFT adopts a discriminative notion, explicitly discriminating good from “bad” outputs generated by the base model to be finetuned. We formalize this approach by introducing a discriminative probabilistic framework that models the discriminative likelihood of an answer among all possible outputs for a given input. This stands in stark contrast to SFT, which uses a generative probabilistic framework to model only the generative likelihood of individual tokens in the answers. To implement this framework, we develop efficient algorithms to optimize the discriminative likelihood of good answers, ensuring both scalability and practicality. Due to its strong discriminative capability, DFT delivers competitive or superior performance compared to SFT→PO, mitigating the requirement for human preference data and reward models.
@@ 2 Related Work
Preference Optimization. Pioneering works proposed RL-based PO methods (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). These methods leverage a separate reward model, trained on human-labeled preference data, and optimize the SFT model against it using policy gradient methods such as PPO (Schulman et al., 2017) and REINFORCE (Williams, 1992). Rafailov et al. (2024) proposed direct preference optimization (DPO), which removes the step of training a reward model and directly optimizes a pairwise loss of the policy model on the preference data. Following DPO, many PO methods have been proposed with different loss functions, including R-DPO (Park et al., 2024), CPO (Xu et al., 2024a), IPO (Azar et al., 2024), SimPO (Meng et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), DPO-p (Pal et al., 2024), to name just a few among others (Zhao et al., 2023; Jung et al., 2024).
Several works (Chen et al., 2024b; Yuan et al., 2023a; Song et al., 2024) have considered PO with a list of ranked preference data that may be explicitly labeled with a reward value. Rosset et al. (2024) assumed a general preference function is given that can produce a probability telling one output is preferred over another output given an input. Different from these works, we do not assume any preference data or preference model other than annotated input-output pairs.
Finetuning via Self-play. Training a model on its own self-generated responses has been widely explored in the PO stage. For example, many variants of RLHF (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022; Li et al., 2023; Chan et al., 2024; Ji et al., 2024) use on-policy samples produced by the current policy under optimization. Some studies have exhibited benefits of using self-generated data for PO by reducing the distribution gap between the training data and the current model while fostering exploration of diverse response spaces (Xu et al., 2024b; Tajwar et al., 2024; Tang et al., 2024). Moreover, leveraging synthetic data has proven essential for iterative (online) algorithmic improvement of these methods (Xu et al., 2023b; Guo et al., 2024; Yuan et al., 2024; Chen et al., 2024a; Dong et al., 2024). A more closely related work is SPIN (Chen et al., 2024c), which uses a similar preference optimization objective as DPO but with data generated by the model to be finetuned as the losing responses. Although we also use self-generated data from the base model to be finetuned as our negative data, our formulation is derived from a discriminative learning framework, making our approach aided by advanced optimization better than the pairwise loss function used in SPIN and other pairwise preference optimization objectives (cf. Section 6).
@@ 4 DFT: Discriminative Finetuning
To address this issue, the current mainstream approach is to finetune the model further using PO on human preference data. If humans label the two answers such that $\mathbf{y}\succ\mathbf{y}^{\prime}$, the model might be able to push the likelihood of $\mathbf{y}^{\prime}$ given $\mathbf{x}$ smaller than that of $\mathbf{y}$ given $\mathbf{x}$. As a result, the good answer $\mathbf{y}$ will be generated more likely than the bad answer $\mathbf{y}^{\prime}$.
@@
Finally, we note that while both DFT and traditional discriminative classification methods (e.g., logistic regression) use supervised data $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i}),i=1,\ldots,n\}$ and model discriminative likelihood through softmax functions, the differences are: (1) DFT operates over an infinite space of possible text outputs $\mathbf{y}^{\prime}$, whereas traditional classification works with a finite set of class labels; (2) the summation over all possible $\mathbf{y}^{\prime}$ in DFT requires advanced optimization techniques, while summation over class labels in traditional approaches is straightforward. This fundamental distinction necessitates our novel sampling and estimation approach while maintaining the core advantages of discriminative learning principles.
@@ 4.1 The Scoring Function
The above framework is similar to the discriminative probabilistic modeling for self-supervised representation learning (Wang et al., 2025). However, we cannot directly borrow the same idea of discriminative representation learning to design the scoring function. In particular, discriminative representation learning uses an encoder network $e(\mathbf{x})$ to induce an embedding of any input text $\mathbf{x}$, and computes the scoring function by using the cosine similarity between $e(\mathbf{x})$ and $e(\mathbf{y})$. However, this representation model $e(\cdot)$ is of no use for generative tasks of LLMs.
@@ 4.2 The Optimization Algorithm
Although our DFT formulations (2) and (3) are nearly identical to that of the traditional discriminative probabilistic approach for classification (e.g., logistic regression), the key challenge lies in solving the optimization problem in (3), particularly in handling the second term of $F(\theta)$, where $\mathcal{Y}$ encompasses all possible texts. Indeed, the optimization problem in (3) is an instance of empirical X-risk minimization (Yang, 2022; Yuan et al., 2023b). We address the optimization challenge by employing advanced optimization techniques of finite-sum coupled compositional optimization framework (FCCO) (Wang & Yang, 2022). The idea is to write the second term of $F(\theta)$ into the form of $\frac{1}{n}\sum_{i=1}^{n}f(\mathbb{E}_{\zeta}g_{i}(\theta;\zeta))$, where $\zeta$ is some random variable. To this end, we introduce a sampling distribution $P_{i}(\cdot)$, which is specified later. Then we define
@@ 7 Experiments
Setup: We evaluate our proposed DFT framework under two distinct training settings. First, we focus on improving the mathematical reasoning capability of a base LLM by using DFT on the MetaMathQA dataset (Yu et al., 2024), which contains 395K samples generated through bootstrapped mathematical questions with augmented reasoning paths. In this setting, we set $B=4$. Second, we fine-tune a base LLM using DFT on the UltraFeedback (UF) dataset (Cui et al., 2023), comprising 61K samples. UF is originally used as a preference dataset, where each data pairs a winning response $\mathbf{y}_{w}$ and a losing response $\mathbf{y}_{l}$. For SFT and DFT, we regard the winning responses $\mathbf{y}_{w}$ as the ground-truth and discard all losing responses. In this setting, we set $B=2$ and generate $\mathbf{y}^{\prime}$ by adding an adversarial prompt like “You are an unhelpful assistant.” to the input $\mathbf{x}_{i}$ in a chat template (cf. Appendix B.5) and use it as input to the base LLM for generation. For both settings, we use Mistral-7B-v0.1 as our base model. More details of implementation and hyper-parameter tuning are described in Appendix A.
@@ 7.1 Results
Observation 2: DFT variants consistently outperform PO methods on self-play data. In Table 3, we compare DFT(2) with PO approaches using self-play data that is generated by the base model as negative data, including SPIN, SimPO, KTO, ORPO, DPO-p and SimPO-SFT, where the last one just combines the SimPO loss and the SFT loss similar to Xu et al. (2024a). Comparing with SimPO-SFT allows us to verify the advantage of our objective over SimPO loss. For these baselines, we use the same generated $\mathbf{y}^{\prime}$ as in DFT as their negative data and finetune the same base model. The results in Table 3 show that these PO approaches on self-play data can not improve SFT. This is different from the observation in Chen et al. (2024c); Xu et al. (2024a), as their experiments are for finetuning an SFT model. Comparing DFT(2) with SPIN and SimPO can justify the effectiveness of our objectives and the optimization algorithm.
@@
Table 2: Comparison between DFT, SFT, and PO methods on self-play data in the second training setting. All methods use the same UF winning responses as positive examples and the same generated outputs from the base model as negative examples, ensuring a fair comparison.
@@
Table 3: Comparison between DFT and SFT→PO approaches on preference data in the second training setting. DFT use only the UF winning responses, while SFT→PO methods use explicit preference pairs.
@@
Table 4: Comparison of DFT with SFT→PO methods without human preference data, where all methods use the same generated outputs from the base model as negative examples.
Method MMLU TruthfulQA HellaSwag Winogrande GSM8k ARC IFEval Avg.
DFT 61.69 52.23 83.95 78.37 48.22 64.25 51.20 62.84
DFT2 61.66 54.14 83.20 77.82 45.49 64.42 51.20 62.56
[…]
@@
AlpacaEval2 Results. Finally, we briefly discuss the GPT4-as-a-judge evaluation on instruction following by reporting the length-controlled winning rate (LC) on AlpacaEval2. The comparison between DFT(2) with SFT and PO-based methods on self-play data is shown in Figure 2(a), which demonstrates DFT(2) outperforms these baselines. The comparison between DFT(2) and SFT→PO approaches on preference data is shown in Appendix B.1, which shows that DFT(2) is competitive with some PO approaches using the preference data, such as KTO, but worse than SimPO and DPO. However, DFT(2) have much shorter lengths for the outputs with an average length of 1359. In contrast, KTO, DPO, SimPO have average lengths of 1449, 1477, 1868, respectively. The GPT4-as-a-judge evaluation tends to favor outputs that are longer. Nevertheless, DFT has competitive if not better performance on verifiable instruction following benchmark IFEval (cf. Table 3).
@@ 7.2 Ablation Studies
We present more results to illustrate the effectiveness of DFT compared with SFT and PO-based objectives, the advantage of our optimization using moving-average estimators $u$, and the effect of the number of generated samples $B$ in each iteration.
Training Curves for the Log-likelihood. Figure 3(a) and Figure 3(b) illustrate the learning dynamics of different methods by tracking the log-likelihood of positive and generated negative examples during training. We compare DFT with SFT and SimPO as in the Table 3. For positive examples (Figure 3(a)), DFT maintains a trajectory similar to SFT, while SimPO shows a decreasing trend, which means the objective of SimPO is not effective if using the self-generated negative examples for PO. For self-generated examples (Figure 3(b)), DFT successfully decreases the log-likelihood, demonstrating its effectiveness in distinguishing between positive and negative examples.
(a) Log-likelihoods of positives
(b) Log-likelihoods of negatives
Figure 3: (a) Log-likelihoods of positive examples during training for different methods. (b) Log-likelihoods of negative examples during training for different methods.
Comparison with SFT→PO without Preference data Table 4 presents results comparing DFT methods with SFT→PO without human preference data, where all methods use the same generated outputs from the base model as negative examples. We use the same UF→UF pipeline as in Table 3, while the only difference is that we replace the losing responses with the same generated responses as DFT in the PO stage of SFT→PO methods. These results demonstrate that the single stage training of DFT(2) is more effective than the two-stage training of SFT→PO methods for using the self-generated data.
@@ 8 Conclusion
Acknowledgments
S. Guo, V. Balmaseda, and T. Yang were partially supported by the NSF SCH grant 2306572.
@@ Appendix A Implementation Details
A.1 Similarity Score
In all experiments, we use the unnormalized generative score $s_{\theta}(\mathbf{y},\mathbf{x})=\log P_{g}(\mathbf{y}|\mathbf{x})$ for DFT, and use the normalized generative score $s_{\theta}(\mathbf{y},\mathbf{x})=\frac{1}{|\mathbf{y}|}\log P_{g}(\mathbf{y}|\mathbf{x})$ for DFT2.
A.2 Hyper-parameters Tuning
@@ A.3 Implementation of SFT→PO
A.4 Training Costs
We compare the training efficiency of different methods, with results shown in Figure 4. For DFT methods, the time is for training of 2 epochs. For SFT→SimPO and SFT→DPO, training time is split into two phases: an initial SFT phase of 2 epochs followed by a preference optimization phase of 1 epoch. Despite the higher computational cost compared to SFT alone, DFT2 offers comparable efficiency to the two-stage SFT→DPO pipeline. All experiments were conducted on 4×A100 80G GPUs. Generation time for negative samples (approximately 1.33 hours for both DFT methods) is not included in this comparison, as it can be performed offline as a preprocessing step.
Figure 4: Comparison of total training time (in hours) for DFT methods and SFT→PO methods on the UltraFeedback dataset.
@@ A.4 Training Costs
A.5 Benchmark Details
Table 7 provides detailed information about the evaluation protocol used for each benchmark in the second training setting.
Table 7: Benchmark evaluation details including number of shots, metrics, and use of chat templates.
Benchmark Shot(s) Metric Chat Template
GSM8k 5 strict-match
ARC 25 acc_norm
HellaSwag 10 acc_norm
TruthfulQA 0 acc
MMLU 5 acc
Winogrande 5 acc
IFEval 0 prompt_level_strict